Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

defective die

Expert replies
by ruplun » Fri Jul 29, 2011 5:20 am
In a defective six-sided dice the probability of getting an odd number is twice the probability of getting an even number.What is the probability of getting 5 in a single throw?

a. 1/18
b. 1/9
c. 2/9
d. 1/2
e. 2/3
Join the discussion
Source: — Problem Solving |

by naveen451 » Fri Jul 29, 2011 9:05 am
IMO C
Join the discussion

by ruplun » Fri Jul 29, 2011 9:05 pm
plz explain the logic
Join the discussion

by ntamhane » Fri Jul 29, 2011 9:29 pm
What's the OA?

My Logic:
In a non defective or unbiased dice the Probability of getting a odd number or a even number is 1/2
The question suggests it is a six sided defective dice with the probability of getting a odd number is 2 times getting the probability of an even number.
Hence the numbers on the dice will be
odd odd odd
even even odd
In this case the Probability of getting the odd number is 4/6 = 2/3
Probability of getting the even number is = 2/6 = 1/3
which satisfies the condition "probability of getting a odd number is 2 times getting the probability of an even number"
Now the extra odd number could be 5 or any other odd number.
i.e
1 3 5
2 4 5
OR
1 3 5
2 4 3(or any other)
hence the probability of getting a 5 in a single throw is 2/6 or 1/6
[spoiler]2/6 + 1/6 = 3/6 = 1/2[/spoiler]
Join the discussion

by sumasajja » Sat Jul 30, 2011 3:27 am
actually given here that the probability of throwing odd numbers is 2
so in a dice for numbers 1 2 3 4 5 6
the sample space is 2+1+2+1+2+1=9
and for getting 5 on dice the defective probability is 2
so probability of getting '5' is
=no.of outcomes /sample space
=2/9
Join the discussion

by Tani » Sat Jul 30, 2011 4:33 am
ntamhane -

Your analysis starts correctly. The probability of getting an odd number is indeed 2/3. But, there are three odd numbers so the probability of getting any specific odd number is 1/3 of the probability of getting any odd number, so (2/3)/3 = 2/9
Tani Wolff
Join the discussion

by ruplun » Sat Jul 30, 2011 8:41 am
prob of getting any odd number is 2/3 how come ? i find its 1/3
Join the discussion

by Tani » Sat Jul 30, 2011 7:43 pm
The probability of getting an odd number is twice the probability of getting an even number. Together they must total 1. So the probability of odd is 2/3 and the probability of even is 1/3.
Tani Wolff
Join the discussion